THEORETICAL AND EXPERIMENTAL STUDY OF THE IR SPECTRUM OF THE BrHI AND BrDI AN-IONS

\underline{TIAO XIE}, ALEXEY KALEDIN, JOEL M. BOWMAN, Department of Chemistry and Cherry L. Emerson Center for Scientific Computations, Emory University, Atlanta, GA 30322; STUART CARTER, Department of Chemistry, University of Reading, Reading RG6 2AD, England; MATT NEE, DANIEL M. NEUMARK, ANDREAS OSTERWALDER, KNUT ASMIS, Department of Chemistry,University of California, Berkeley, CA 94720.

A new potential energy surface and dipole moment surface were constructed for BrHI^{−} by a 3-D spline interpolation of ab initio data at the MRCI/aVQZ level of theory and basis set, with ECP’s for the halogens. ”Exact” vibrational calculations were performed on this potential energy surface for BrH(D)I^{−} using two very different codes. The asymmetric stretch and bending modes are found to be strongly coupled, and are shown to be in close 1:2 Fermi resonance. This is verified by a simple de-perturbation of the wavefunctions. The IR transition probabilities were also obtained by integrating dipole moment matrix elements over both vibrational and rotational coordinates. Excellent agreement was found with the recently measured IR spectrum of BrH(D)I^{−} using the Ar-messenger technique by Neumark and co-workers.